OpenStudy (anonymous):

evaluate cos 220 cos40+sin220sin40

OpenStudy (anonymous):

This question came up earlier, I think spanishkb31 is in your class...

OpenStudy (anonymous):

okay, cos(-40) = cos(40) and cos(a+b) = cos(a)cos(b) - sin(a)sin(b)...

OpenStudy (anonymous):

Nice memory :o

OpenStudy (anonymous):

so what is (a+b)?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you all set?

OpenStudy (anonymous):

no lost

OpenStudy (anonymous):

whats the step

OpenStudy (anonymous):

well, if you replace cos(40) with cos(-40), you get: cos 220 cos(-40) - sin220sin(-40) which is the sum of angles formula for cosine cos(a+b) = cosa cosb - sina sinb so, what is (a+b) ?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

whats the answer